75g KMnO4=. Molecules KMnO4


1
Expert's answer
2021-03-15T09:28:00-0400

Solution:

Moles of KMnO4 = Mass of KMnO4 / Molar mass of KMnO4

The molar mass of KMnO4 is 158.034 g mol-1.

Hence,

Moles of KMnO4 = (75 g) / (158.034 g mol-1) = 0.4746 mol KMnO4


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

Number of KMnO4 molecules = (0.4746 mol) × (6.022×1023 molecules / 1 mol) = 2.858×1023 molecules

Number of KMnO4 molecules = 2.858×1023 molecules


Answer: 2.858×1023 KMnO4 molecules.